# Star formation beyond galaxies: widespread in-situ formation of intra-cluster stars in TNG50

### v1

@@ 3 Results
It is also important to note that at the current integrated star formation rates shown in the intra-cluster region of our simulated clusters, SFR 1 /yr (see Fig. 3), it would take Gyr to form the total mass of the in-situ ICL stars listed in Table 1. This diffuse mode of star formation must therefore have also been present at earlier times, and with higher intensity. We confirm this by showing the distribution of stellar mass formed in-situ in a given time-bin in Fig. 7. This is calculated by adding all the stellar mass of the in situ-labeled stars in the ICL with a given age divided by the size of the time-bin to provide an average star formation rate. The blue shaded histograms indicate that the in-situ intra-cluster star-formation is a rather continuous phenomena over time and was, on average, stronger around - (i.e., cosmic times - Gyr). As an illustration, we show in Fig. 8 the predicted SFR maps for our highest mass group at three different redshifts: . In all cases, extended star forming areas appear out to almost the virial radius of the cluster progenitor.
@@
The color coding of the gas in Fig. 6 indicates that, at , the expected H surface brightness for the intra-cluster gas is rather modest, – , which could partially explain why this phenomena has so far escaped detection. For comparison, the most extended and deepest surveys available, such as VESTIGE, have a sensitivity at 3 arcsec angular resolution (Boselli et al., 2018a). On the other hand, there are claims of detection of runaway cooling of the hot gas in high- clusters leading to star formation events with SFR 900 /yr offset from the main host galaxy (Hlavacek-Larrondo et al., 2020) that might be comparable to some of the in-situ formed stars in our simulations.
Finally, we must also consider the possibility that these in-situ formed ICL stars are a numerical effect born out of specific choices made for the baryonic treatment in the simulation. The fact that most star-forming clouds sit near the density threshold for star formation suggests that increasing the value of that parameter might significantly lower the occurrence of stars born in the in-situ ICL component, but more rigorous tests must be performed to quantify this effect.
@@ 4 Summary and Conclusions
We use the cosmological hydrodynamical simulation TNG50 to study the formation of in-situ ICL in Virgo-like systems. We select the 3 most massive clusters in the box with virial masses -. We find that a significant fraction, - of the present-day mass in the ICL, was formed in-situ. When analyzed in detail, and contrary to what is commonly believed, the majority of these in-situ stars are born directly into the diffuse light component at hundreds of kiloparsecs from the main host and directly from the CCM gas, rather than brought in and dispersed during merger events. This diffuse star-formation proceeds in cloudlets that condense along filamentary-like structures that are aligned with the neutral gas and dark matter distributions and are not clearly associated with ram-pressure stripping of gas in infalling galaxies.
At the expected H surface brightness due to this diffuse mode of star formation reaches , just below the detection limits of the ongoing surveys in Virgo when smoothed on scales of arcsec. In the simulated clusters, star formation maps at higher redshift suggest that extended star formation is present throughout the history of the cluster, making it an interesting target for discovery with facilities such as JWST.
One interesting application of our results is that orphan supernova events would be expected to occur in relative isolation in the ICL. While such events have not yet been reported, with a predicted average SFR of /yr, we expect only 1 SN per yr in a Virgo-like system222We have assumed a power-law mass function following Salpeter (1955), for the mass range of – with stars more massive than exploding as SN, which results in of the total stellar mass formed able to generate a SN event. Assuming a mean SN mass , we anticipate the occurrence of 1 SN in () years., suggesting that a concurrent survey of hundreds of clusters might be needed for detecting such rare events.
If ongoing star formation in the intra-cluster region of observed clusters is really lacking, then this would signal shortcomings in the treatment of star formation in simulations like TNG50, especially in low-density regions. The formation of stars in these extended and diffuse structures is possible because of the condensation of denser clouds embedded in the more diffuse CCM, as reported by Nelson et al. (2020). Numerical techniques such as that used in Arepo have been shown to lead to less artificial formation of clouds compared to other techniques such as SPH (Agertz et al., 2009; Torrey et al., 2012), but it is possible that other numerical effects are still at play. We emphasize that further studies of the effect of the star formation threshold in models such as those of TNG would be desirable to determine their role in the formation of stars in such unexpected conditions.
@@ Acknowledgements
We gratefully acknowledge the contributions of several individuals who supported this research. NA and LVS would like to acknowledge the financial support received from the NASA ATP-80NSSC20K0566, NSF-CAREER-1945310 and NSF-AST-2107993 grants. NA would like to acknowledge the support provided by the UCR-Carnegie Graduate Fellowship.
@@
This paper is based on halo catalogs and merger trees from the IllustrisTNG Project (Nelson et al., 2019). These data are publicly available at https://www.tng-project.org/.
Appendix A Effect of resolution
To assess the impact of resolution on the star formation rates found within the ICL regions in our studied groups in TNG50, we turn our attention to the results obtained from TNG100. TNG100-1, part of the same suite of cosmological hydrodynamical simulations, represents a larger baryonic run with a volume of and resolution elements (lower resolution compared to TNG50, which features resolution elements). In this run, the baryon and dark matter particle masses are and , respectively.
For the comparison, we exclusively consider groups with virial masses similar to those examined in our study (), resulting in a subset of 26 groups in TNG100.
In Fig. 9, we juxtapose the star formation rates in TNG50 and TNG100 for both central and ICL regions. The results reveal no discernible difference in the ICL star formation rates among various groups between the two simulations (represented by navy squares for TNG50 results and light blue squares for TNG100 results), aside from a larger scatter in the TNG100 results, primarily due to the larger sample size in the more extensive box of TNG100).

### v2

@@ 3 Results
At z=0, the velocity distributions of both the in-situ and accreted stellar components are found to be remarkably similar, suggesting significant virialization within the cluster. This hints at the possibility that the in-situ component must also be relatively old, comparable in age to the accreted stars. We also note that at the current integrated star formation rates shown in the intra-cluster region of our simulated clusters, SFR 1 /yr (see Fig. 3), it would take Gyr to form the total mass of the in-situ ICL stars listed in Table 1. This diffuse mode of star formation must therefore have also been present at earlier times, and with higher intensity. We confirm this by showing the distribution of stellar mass formed in-situ in a given time-bin in Fig. 7. This is calculated by adding all the stellar mass of the in situ-labeled stars in the ICL with a given age divided by the size of the time-bin to provide an average star formation rate. The blue shaded histograms indicate that the in-situ intra-cluster star-formation is a rather continuous phenomenon over time and was, on average, stronger around - (i.e., cosmic times - Gyr). As an illustration, we show in Fig. 8 the predicted SFR maps for our highest mass group at three different redshifts: . In all cases, extended star forming areas appear out to almost the virial radius of the cluster progenitor. An examination of the birthplaces of the in-situ stars in the ICL reveals that a fraction of these stars may have originated in the central galaxy and were subsequently ejected due to merger events. This phenomenon is especially noticeable in one of our groups, which experienced a recent major merger in its formation history. However, for the rest of our sample, we estimate that the majority (over 60%) of the in-situ ICL stars formed at kpc or , relative to the BCG or its main progenitor at the time of their formation.
These stars are forming from the cool gas present in the ICM of the clusters, as shown in the star-forming gas distributions at higher redshifts in Fig. 8. The existence of cool gas in filamentary structures around these clusters at higher redshift () has also been recently reported in the study by Rohr et al. (2024), which examines more massive clusters (with masses ranging from to ) from the TNG-Cluster simulation suite (Nelson et al., 2024). This suite employs the same baryonic physics as TNG50 but at a lower resolution. Our measured SFRs in the ICM for the mass range of our three systems are roughly in agreement with their reported SFRs at where the mass range of their more massive clusters overlaps with our systems.
@@
The color coding of the gas in Fig. 6 indicates that, at , the expected Hα surface brightness for the intra-cluster gas is rather modest, – , which could partially explain why this phenomena has so far escaped detection2222These quoted values for brightness are derived from Fig. 6, where we binned the star-forming gas at similar spatial resolution as VESTIGE (3 arcsec) and measured the flux within each ’pixel“. If we analyze the results and compute the median and dispersion, we obtain a median value of and a dispersion ranging from – .. For comparison, the most extended and deepest surveys available, such as VESTIGE, have a sensitivity at 3 arcsec angular resolution (Boselli et al., 2018a)3333The stacking of pixels to increase image sensitivity of the images should be limited to reasonable scales. The 3 arcsec scale ( 0.24 kpc at the distance of the Virgo cluster) mentioned in the VESTIGE paper is optimized for detecting structures associated with the tails of stripped gas in galaxies. While stacking pixels on 100 kpc scales could theoretically increase sensitivity as (where N is the number of pixels), this assumes a perfectly flat background, which is rarely the case. Issues such as large-scale structures from flat fielding imperfections and correlated noise due to bright stars and unwanted gradients from non-uniform illumination can affect the results (see Fig. 4 of Boselli et al. (2018a), VESTIGE paper I). Moreover, narrow-band filters centered on the rest-frame line are also sensitive to emission from Galactic cirrus. Hence the contribution of the Galactic cirrus in the direction of the Virgo cluster can complicate measurements (see Fig. 16 of the same VESTIGE I paper). While IFU spectroscopy could help confirm these features, technical limitations exist. For instance, MUSE/VLT can achieve on 1 arcsec scales in 1h integration, but its 1 arcmin FoV is insufficient for mapping large areas like the Virgo cluster (100 sq.deg), necessitating pre-selection of observation targets.. On the other hand, there are claims of detection of runaway cooling of the hot gas in high- clusters leading to star formation events with SFR 900 /yr offset from the main host galaxy by kpc for the coolest gas and kpc for the X-ray peak (Hlavacek-Larrondo et al., 2020) that might be comparable to some of the in-situ formed stars in our simulations.
Finally, we must also consider the possibility that these in-situ formed ICL stars are a numerical effect born out of specific choices made for the baryonic treatment in the simulation. While the multi-phase nature of the gas in the outskirts of groups and clusters has been reported before by several simulation groups (Nelson et al., 2018b; Butsky et al., 2019; Oppenheimer et al., 2021), whether warm/cold gas cells in the ICM can form stars is less robustly predicted by numerical codes. The fact that most star-forming clouds sit near the density threshold for star formation suggests that increasing the value of that parameter might significantly lower the occurrence of stars born in the in-situ ICL component, but more rigorous tests must be performed to quantify this effect. More detailed baryonic prescriptions that explicitly resolve the multiphase structure of the gas in the ISM of galaxies advocate for larger density thresholds for star formation, - , more in line with the typical densities of molecular clouds with active star formation (Hopkins et al., 2011; Agertz et al., 2013; Dutton et al., 2019). Our study points out that baryonic treatments that are able to successfully reproduce many of the properties of galaxy populations, might not be as well attuned to perform well in environments where the typical conditions of the gas differs from that in the ISM of galaxies, for example in the ICM. As shown in Fig. 8, the potential limitations of these treatments might worsen at intermediate to high-redshifts, where the associated higher densities translate into larger numbers of gas cells being eligible to become star-forming in prescriptions based on a low density threshold for star formation. In other words, while baryonic treatments such as the one employed in the TNG model, do predict the presence of overdense pockets of gas within the ICM, it is insufficient to uniquely determine which fraction of that gas is dense and cold enough to form stars, a limitation that may scale with redshift as gas is more abundant and dense at high-z. As more data become available for low- and high-z groups and clusters, their imaging in wavelengths where young stars are mapped might help understand the applicability of such star formation recipes to these environments.
@@ 4 Summary and Conclusions
We use the cosmological hydrodynamical simulation TNG50 to study the formation of in-situ ICL in Virgo-like systems. We select the 3 most massive clusters in the box with virial masses -. We find that a significant fraction, - of the present-day mass in the ICL, was formed in-situ. When analyzed in detail, and contrary to what is commonly believed, the majority of these in-situ stars are born directly into the diffuse light component at hundreds of kiloparsecs from the main host and directly from the ICM gas, rather than brought in and dispersed during merger events. This diffuse star-formation proceeds in cloudlets that condense along filamentary-like structures that are aligned with the neutral gas and dark matter distributions and are not clearly associated with ram-pressure stripping of gas in infalling galaxies.
At the expected Hα surface brightness due to this diffuse mode of star formation reaches , just below the detection limits of the ongoing surveys in Virgo when smoothed on scales of arcsec. In the simulated clusters, star formation maps at higher redshift suggest that extended star formation is present throughout the history of the cluster, making it an interesting target for discovery with facilities such as JWST.
One interesting application of our results is that orphan core-collapse supernova events would be expected to occur in relative isolation in the ICL. While such events have not yet been reported, with a predicted average SFR of /yr, we expect only 1 SN per yr in a Virgo-like system4444We have assumed a power-law mass function following Salpeter (1955), for the mass range of – with stars more massive than exploding as SN, which results in of the total stellar mass formed able to generate a SN event. Assuming a mean SN mass , we anticipate the occurrence of 1 SN in () years., suggesting that a concurrent survey of hundreds of clusters might be needed for detecting such rare events.
If ongoing star formation in the intra-cluster region of observed clusters is really lacking, then this would signal shortcomings in the treatment of star formation in simulations like TNG50, especially in low-density regions. The formation of stars in these extended and diffuse structures is possible because of the condensation of denser clouds embedded in the more diffuse ICM, as reported by Nelson et al. (2020). Numerical techniques such as that used in Arepo have been shown to lead to less artificial formation of clouds compared to other techniques such as SPH (Agertz et al., 2009; Torrey et al., 2012), but it is possible that other numerical effects are still at play. We emphasize that further studies of the effect of the star formation threshold in models such as those of TNG would be desirable to determine their role in the formation of stars in such unexpected conditions.
@@ Acknowledgements
We gratefully acknowledge the contributions of several individuals who supported this research. Special thanks to our anonymous referee whose valuable suggestions greatly improved this work. The authors would also like to thank Eric Rohr, Annalisa Pillepich, Dylan Nelson, Lars Hernquist and Chris Mihos for insightful discussions that helped shape this paper. NA and LVS would like to acknowledge the financial support received from the NASA ATP-80NSSC20K0566, NSF-CAREER-1945310 and NSF-AST-2107993 grants. NA would like to acknowledge the support provided by the UCR-Carnegie Graduate Fellowship.
@@
This paper is based on halo catalogs and merger trees from the IllustrisTNG Project (Nelson et al., 2019a). These data are publicly available at https://www.tng-project.org/.
Appendix A Physical conditions for star formation
In this section we provide additional information on the gas density of the star-forming gas. We show in Fig. 9 the normalized distribution of gas density of star-forming gas in different environments: ICM (blue), satellites(gray) and central cluster galaxies (orange) in our clusters. The classification of cells as either belonging to the ICM or to satellites has been done consistent with the classification of stars, following membership of cells as assigned by Subfind and “cleaning” the ICM of particles close in distance to surviving satellites (see Sec. 2.1). Fig. 9 shows that the star formation in the ICM occurs with densities biased slightly low with respect to the star formation in the ISM of galaxies, but there is considerable overlap in the distributions.
Appendix B Effect of resolution
To assess the impact of resolution on the star formation rates found within the ICL regions in our studied groups in TNG50, we turn our attention to the results obtained from TNG100 and TNG50-2. TNG100-1, part of the same suite of cosmological hydrodynamical simulations, represents a larger baryonic run with a volume of and resolution elements (lower resolution compared to TNG50, which features resolution elements). In this run, the baryon and dark matter particle masses are and , respectively. TNG50-2 has the exact same volume as TNG50 but simulated at the resolution of TNG100.
For the comparison, we exclusively consider groups with virial masses similar to those examined in our study (), resulting in a subset of 26 groups in TNG100 and 3 groups in TNG50-2.
In Fig. 10, we juxtapose the star formation rates in TNG50 and TNG100 for both central and ICL regions.
The results of TNG50 seem within the scatter of those for the larger sample of TNG100 groups, perhaps slightly biased to the upper half of the SFR distribution. For a more direct comparison, we also show the 3 most massive groups in the TNG50-2 run, close in resolution to TNG100. These comparisons suggests a weak dependence of the results with numerical resolution which is in line with the resolution trends found in the formation of clouds in Nelson et al. (2020): better resolved runs seem more prone to form clouds and also will result in a larger chance to have star-forming gas in them.